OpenStudy (anonymous):

Medal & fan. Determine the solution set of (3x + 1)^2 - 100 = 0.

OpenStudy (anonymous):

3x + 1 = 100 right?

OpenStudy (freckles):

Hey you might want to check your previous question. I think you were under the wrong impression what the answers were. well first add 100 on both sides (3x+1)^2=100 then take square root of both sides you nee to take square root of 10 and you should have a second equation 3x+1=10 or 3x+1=-10

OpenStudy (freckles):

(x+2)^2=36 gives the equations x+2=6 or x+2=-6 subtracting 2 on both sides gives x=6-2 or x=-6-2 which gives us x=4 or x=-8

OpenStudy (freckles):

8 is definitely not a solution because (8+2)^2=36 is totally false

OpenStudy (anonymous):

so its not (4, -8) @freckles

OpenStudy (freckles):

i just said it is x=4 or x=-8

OpenStudy (freckles):

I was disagreeing with ever gave you 8 as an answer

OpenStudy (anonymous):

oh myb i see now

OpenStudy (freckles):

as (8+2)^2=36 is false (-8+2)^2=36 is true (4+2)^2=36 is also true

OpenStudy (freckles):

anyways can you solve the two linear equations above

OpenStudy (freckles):

3x+1=10 or 3x+1=-10